OpenStudy (anonymous):

Explain how you would solve the equation -7 + x = 12 using the properties of equality.

OpenStudy (alfie):

Add seven to both sides, so you get.. 7 - 7 + x = 12+7 Now you get.. x= 19

OpenStudy (anonymous):

Thanks